# Towards the New XAI: A Hypothesis-Driven Approach to Decision Support Using Evidence

### v3

@@ Abstract
Prior research on AI-assisted human decision-making has explored several different explainable AI (XAI) approaches. A recent paper has proposed a paradigm shift calling for hypothesis-driven XAI through a conceptual framework called evaluative AI that gives people evidence that supports or refutes hypotheses without necessarily giving a decision-aid recommendation. In this paper, we describe and evaluate an approach for hypothesis-driven XAI based on the Weight of Evidence (WoE) framework, which generates both positive and negative evidence for a given hypothesis. Through human behavioural experiments, we show that our hypothesis-driven approach increases decision accuracy and reduces reliance compared to a recommendation-driven approach and an AI-explanation-only baseline, but with a small increase in under-reliance compared to the recommendation-driven approach. Further, we show that participants used our hypothesis-driven approach in a materially different way to the two baselines.
@@ 1 Introduction
Research has shown that AI recommendations, even when accompanied with explanations, are not always helpful in supporting decision-making [4, 39, 6, 15]. The direct causes of this are under-reliance and over-reliance [38]. With under-reliance, decision-makers reject AI recommendations, even when they may be correct. Alternatively, decision-makers may overly rely on AI recommendations, hence be led to errors when the AI is incorrect. In either case, they tend to fixate on a particular hypothesis without sufficiently considering others [28]. Approaches such as cognitive forcing, based on ideas from human psychology, have been proposed to address limitations of the AI recommendation approach [6], with recent work indicting that withholding AI model the recommendations, at least for a short time, while still providing the user with an explanation of that recommendation, can be helpful [15]. Recently, Miller [28] proposed a so-called hypothesis-driven decision-making paradigm called evaluative AI. The main aim of this paradigm is to focus the decision loop on the human decision maker, providing them with the right evidence to support their own intuitions, rather than focusing the decision loop on machine recommendations. This paradigm offers a promising direction in building better decision support in explainable AI (XAI) research by focusing on human decision makers considering multiple possible hypotheses.
In this paper, we describe and evaluate an approach for building a hypothesis-driven decision-making model that uses the Weight of Evidence (WoE) framework [2]. To the best of our knowledge, this is the first work that compares empirically and in a controlled manner the hypothesis-driven approach [28] with two other popular decision-making approaches (recommendation-driven and AI-explanation-only [15]). Our contributions are:
@@
  • The Evidence-Informed Hypothesis-Driven Decision-Making model, building on the Weight of Evidence (WoE) framework to the hypothesis-driven approach;
  • Two human behavioural experiments comparing our hypothesis-driven approach with two common decision-aid approaches: (1) the standard model recommendation with explanation; and (2) a form of cognitive forcing that provides only AI explanations [15]. The results show that hypothesis-driven approach improved decision accuracy and reduces over-reliance compared to standard recommendation-driven approaches, at the cost of a slight increase in under-reliance. Furthermore, the hypothesis-driven approach reduces under-reliance significantly compared to the AI-explanation-only approach. Our qualitative analysis identifies some limitations and challenges in the three approaches, and shows that participants used the hypothesis-driven approach in a materially different way than the recommendation-driven or AI-explanation-only conditions, with participants focusing more on the evidence than on their own background knowledge.
@@ 2 Background and Related Work
In this section, we give a brief overview of related work on common AI-assisted decision-making paradigms. We then highlight key literature on applying explainable AI in supporting decision-making.
2.1 AI-Assisted Decision-Making Paradigms
@@
Human-first Workflow
Human-first decision workflow has been shown to help reduce the over-reliance on erroneous AI recommendations [6, 14]. However, experts may interact with decision-making systems differently from laypeople (crowdworkers). For example, Fogliato et al. [14], Gaube et al. [16] run studies with radiologists who were the experts. Their task is to review patients’ X-ray images. The studies conclude that in human-first workflow with expert participants, they are less likely to leverage AI advice even though the AI is more accurate. In fact, this is referred to algorithm aversion [12] or under-reliance.
A human-first paradigm called cognitive forcing, based on earlier ideas in psychology for interventions that elicit human thinking at decision-making time [23], has been proposed as a way to improve users’ engagement and also increase their learning when interacting with the AI [6, 15]. Four cognitive forcing designs have been introduced: (1) On demand: Participants can only see the AI recommendation when they request it; (2) Update: Participants first made a decision without seeing the AI recommendation. Then, they were shown the AI prediction and could update their decision later; (3) Wait: Participants had to wait for 30 seconds before the AI decision was shown; and (4) Only AI explanation: Providing just the AI explanation and no AI recommendation, on the basis that this may help people process the AI explanation more carefully and therefore, improve their knowledge and make better decisions [15]. Importantly, cognitive forcing has been shown to reduce over-reliance compared to the standard AI suggestion approach, although that study had a limitation in that the AI prediction was always correct. There are also some recognised trade-offs of cognitive forcing designs: more time-consuming [15], and less trust [6].
Evaluative AI (Hypothesis-driven) Paradigm
Miller [28] argues that AI-assisted decision support is on the cusp of a paradigm shift. This shift is away from the idea of human-first or AI-first, and into a framework he calls evaluative AI. The key insight of evaluative AI is to not necessarily provide a recommendation, and instead to support the human cognitive decision-making process by providing evidence for or against the particular hypothesis that a human decision maker is considering. Miller argues that this would help to prevent over- and under-reliance, and would help the decision maker to retain their internal locus on control [32].
@@ 2.2 Explainable AI (XAI) in Decision Support
In this section, we review some AI-first explainable AI approaches that have been used to provide explanations for the AI recommendation based on feature analysis. We will also discuss some popular evidence-based explanations that have been used in supporting decision-making.
Human decision-making reliance on AI support
@@
There is no straightforward position regarding when humans are well calibrated to accept AI-generated advice [38]. Overall, study participants appear more likely to accept an AI’s recommendation when provided with explanations, regardless of the model’s correctness [4, 19, 36]. Explanations can increase the accuracy of the human-AI team when the AI is correct, but decrease it when it is wrong, resulting in over-reliance on the AI’s recommendations. Arguably, this is because the current explanation styles do not provide details of the underlying rationale of the AI model behaviour [36]. We therefore should be careful when selecting the explanation type as it can have a significant effect on whether users decide to rely on them  [7, 26].
Evidence-based Explanations
@@
In this paper, we will generate evidence-based explanations, which are similar to feature importance explanations. The main difference is that Weight of Evidence uses log likelihoods and log odds ratios to generate explanations, whereas LIME [31] and SHAP [27] find feature importance by modifying the predictive posterior probability in various ways.
@@
Evidence-based explanations have been applied to support decision-making and debug models in several prior research [20, 21, 29] other than from Alvarez Melis et al. [2]. A closely related work to Alvarez Melis et al. [2] is from Poulin et al. [29]. Poulin et al. [29] propose a framework called ExplainD that uses additive evidence. The framework also measures the weight of evidence using a Naive Bayes classifier along with highlighting the negative and positive evidence for a decision. However, the problem being considered is a binary classification. Furthermore, there is still room for improvement by conducting experiments to evaluate the framework. Kulesza et al. [20, 21] introduce EluciDebug in email classification using Multinomial Naive Bayes classifier (MNB). The EluciDebug prototype provides an interface that includes important words and the folder size that both contribute to the email classification. However, the prototype did not specifically give positive and negative evidence in decision-making situations.
3 Evidence-Informed Hypothesis-Driven Decision-Making Model
We define the evidence-informed hypothesis-driven decision-making model by implementing the evaluative AI (hypothesis-driven) paradigm [28] using the WoE model. Specifically, given a classification problem, decision makers explore evidence for and against each hypothesis (i.e. an output class). We allow decision-makers to interact with the model by repeatedly selecting a hypothesis for which they can then see the positive (or negative) evidence.
Evidence Generation

### v4

@@ Abstract
This paper presents a hypothesis-driven approach to improve AI-supported decision-making that is based on the Evaluative AI paradigm - a conceptual framework that proposes providing users with evidence for or against a given hypothesis. We propose an implementation of Evaluative AI by extending the Weight of Evidence framework, leading to hypothesis-driven models that support both tabular and image data. We demonstrate the application of the new decision-support approach in two domains: housing price prediction and skin cancer diagnosis. The findings show promising results in improving human decisions, as well as providing insights on the strengths and weaknesses of different decision-support approaches.
keywords:
@@ 1 Introduction
Research on AI-assisted human decision-making has explored several different explainable AI (XAI) approaches. A common approach is to provide explanations of the AI model’s predictions, which can help users understand why the model made that prediction. This approach is called recommendation-driven AI, where the model provides a recommendation and additional information about that recommendation. Human decision-makers can then decide whether they should follow the recommendation or not. However, research has shown that AI recommendations, even when accompanied with explanations, are not always helpful in supporting decision-making [5, 84, 9]. In recent research, Miller [60] argues that the recommendation-driven approach has two main issues: (1) explaining just the AI recommendation limits the user’s control, as few alternatives are offered if the user disagrees with the AI; (2) the recommendation-driven approach does not align with the cognitive processes of human decision-making. Therefore, challenges remain in building more effective AI-assisted systems in the future.
One challenge is that the recommendation-driven approach can cause algorithm aversion [20] where people do not trust the AI; or worse, over-reliance on the AI system [82]. Importantly, research has shown that providing explanations does not always reduce over-reliance on AI recommendations, compared to only providing AI predictions [5, 9]. More troubling is the fact that adding more information (i.e., adding explanations for the AI recommendation) is not always helpful and sometimes can lead to worse performance and overconfidence in the wrong information [67, 66]. For instance, explanations can result in people viewing the model’s incorrectness as being correct [34] and people being deceived by incorrect explanations regardless of their expertise [61].
Another challenge is that providing recommendations and explanations does not align well with accepted cognitive processes of human decision-making [68, 41, 29, 60]. Specifically, Hoffman et al. [29] argue that abductive reasoning is an appropriate basis for conceptualising explainable AI, as it involves putting forward hypotheses to explain an event. Further, based on Yates and Potworowski [87]’s definition of cardinal decision issues, Miller [60] identifies six criteria for good decision aid, including: (1) help to identify options, (2) help to identify possible outcomes for each option, (3) help to judge which outcomes are most likely, (4) help to identify impacts on stakeholders, (5) help to make trade-offs between options, and (6) help to understand the machine’s decision. The current recommendation-driven approach (i.e., giving AI recommendations and explanations) only satisfies criteria (6).
To address these issues, Miller [60] proposed a paradigm shift called hypothesis-driven decision support using a conceptual framework termed evaluative AI. The new framework aims to provide evidence for or against a given hypothesis. Rather than telling the decision-makers what to do, the hypothesis-driven approach allows a user to have more control of the decision-making process by incorporating their hypotheses and promoting uncertainty awareness [49, 51].
In recent work [51], we proposed a particular evidence-informed hypothesis-driven decision-making model obtained by applying the Weight of Evidence (WoE) framework [59]. The results showed that the hypothesis-driven approach can improve decision accuracy and reduce over-reliance on AI, but with a small increase in under-reliance when compared to the recommendation-driven approach, in a user study of housing price prediction. This paper is an extension of that prior work [51] in two ways. First, we extend the Weight of Evidence (WoE) framework from tabular data to (unstructured) image data by applying concept-based explanations [89, 88] to extract high-level concepts, which represent the features of the image. The original WoE framework had been implemented only for tabular data, while extracting features from images is more complex than processing pre-defined features in tabular data. Second, using a case study of supporting skin cancer diagnosis, we report on a human behavioural experiment that investigated the differences between the two approaches (recommendation-driven and hypothesis-driven). To the best of our knowledge, this is the first work that proposes a general tool for Evaluative AI.
Our contributions are as follows:
@@
  • A hypothesis-driven approach for decision support by providing evidence for or against a given hypothesis, built on the Weight of Evidence (WoE) framework.
  • Extension of the Weight of Evidence (WoE) framework to support both tabular and image data. The latter refers to a Visual Evaluative AI model. We also provide public access to this implementation as a Python package  [^1] https://github.com/thaole25/EvaluativeAI.
  • Evaluating the hypothesis-driven approach compared to the previous decision support models (AI-explanation-only [23] and recommendation-driven) through crowdsourcing experiments with laypeople in the housing price prediction domain. The results show that the hypothesis-driven approach can improve decision accuracy and reduce over-reliance on AI, but with a small increase in under-reliance.
  • Evaluating the hypothesis-driven approach in the skin cancer diagnosis domain. Study participants are individuals with a background in the skin cancer field (e.g., PhD students, postdoctoral researchers, medical doctors and melanographers). The results show that both recommendation-driven and hypothesis-driven have pros and cons, and the hypothesis-driven approach is more preferred by experienced diagnosers. Based on this study, we also propose suggestions for the design of future decision support systems.
The following sections of this paper are organised as follows: Section 2 covers literature review on human cognitive processes when making decisions, different decision-support paradigms and applications of explainable AI in building decision support; Section 3 describes the methodology of the proposed decision support models, including the Weight of Evidence (WoE) model and the concept-based explanation approaches; Section 4 presents the study in housing price prediction; Section 5 details the study in skin cancer diagnosis; and Section 6 discusses the results of the study and suggestions for future research. The study instruments are available in our online supplementary materials.
@@ 2 Background and Related Work
In this section, we first discuss the cognitive processes in human decision-making and the cognitive biases that form the basis for implementing abductive reasoning in decision support [29, 60]. Subsequently, we review different decision support paradigms that inform the experimental designs of our studies. This section concludes with an overview of how explainable AI has been previously applied in decision support, including the reason for choosing the Weight of Evidence (WoE) approach [59] as the foundation for our work.
2.1 Cognitive Processes in Human Decision-Making
A well-known psychology theory is the dual process theory [32], which suggests that humans have two different systems for processing information: (1) System 1 is fast, automatic and intuitive; and (2) System 2 is slow, deliberate and more accurate. Aligning with this theory, there are two different types of decisions that can be made by humans [43]: (1) reflexive decision and (2) multi-attribute decision. A reflexive decision is a simple decision made instantly in a short time. It neither involves many attributes in the input nor demands conscious thoughts. By contrast, a multi-attribute decision is a complicated decision that involves many attributes and numerous alternatives.
We might think that System 2 (slow) is better than System 1 (fast) in decision-making and should aim to avoid System 1. However, Miller [60] argues against this view and suggests that we should use the strengths of both systems in designing decision aids. Specifically, the Naturalistic Decision Making (NDM) community considers intuition as prior experience that can be used to make rapid and accurate decisions without having to evaluate all options [38, 42]. The value of intuition can be overlooked if we conduct experiments with laypeople who have no experience in the task and, therefore, fail to capture the advantages of prior knowledge in decision-making [38].
Then how should we design decision-support systems that can leverage both System 1 and System 2? Klein et al. [39, 40] examine sensemaking from various psychological perspectives, in which sensemaking refers to how people make sense of the world. Klein et al. [40, 41] present a theory of sensemaking known as Data/Frame Theory. The data is the information and observations that we use to reconstruct the frame, which is a generalisation of a hypothesis. We can question the frame and seek new data to adjust the frame. This process is iterative and done using both System 1 and System 2. People make their decisions first by using their prior knowledge (System 1, System 2) and then search for evidence and make deliberate judgements among plausible options (System 2). Along the same lines, Hoffman et al. [29] argue that Peirce’s notion of abductive reasoning [68] best reflects the cognitive processes in the XAI model. Abductive reasoning is the cognitive process when we give hypotheses to explain an occurred event. Based on this foundation, we will comment on the idea of Evaluative AI [60] in Section 2.3.3 and elaborate our method in Section 3.
2.2 Cognitive Biases in Decision-Making
Cognitive biases, introduced by Tversky and Kahneman [79], represent systematic errors in judgment, affecting how individuals perceive input information. These biases can lead to inaccurate and irrational decisions in decision-making contexts. Therefore, understanding the effects of cognitive biases on human-AI decision-making settings, with a particular focus on the application of XAI techniques, is important.
Tversky and Kahneman [79] outlined several heuristics and biases such as availability, representativeness, and anchoring. The availability heuristic refers to the tendency to rely on the information that is easiest to recall in memory during the evaluation process. Representativeness involves evaluating the probability of an event based on how similar it is to a typical case, rather than measuring the true statistical probability of the event. Anchoring refers to the reliance on the first piece of information encountered when making decisions. Tversky and Kahneman [79]’s foundational work has continued to inform much of the literature on cognitive biases such as confirmation bias [37, 64], automation bias [54], framing [80], fixation [41], etc.
The goal of XAI is to enhance the interpretability and transparency of complex AI models, thereby improving human understanding and trust in AI. Despite the benefits of XAI, human cognitive biases can still influence the decision-making process. For instance, Bertrand et al. [6] provided a systematic review of the connection between cognitive biases and XAI. They found that cognitive biases can affect or be affected by XAI in various ways, which were classified into four categories: (1) cognitive biases that affect how XAI methods are designed (e.g., explanatory heuristics); (2) cognitive biases that occurred in user studies (e.g., preference for usability over accuracy); (3) cognitive biases that can be mitigated by XAI (e.g., providing prototypes to mitigate representativeness bias); and (4) cognitive biases that can be worsened by XAI (e.g., confirmation bias can lead to over-reliance on the AI). The recommendation-driven approach is an example of how confirmation bias can lead to over-reliance.
These biases can compromise decision quality even with the help of XAI methods; therefore, de-biasing strategies are needed to reduce the impact of human cognitive biases on decision-making. The idea is to provide explanations for not only the AI’s prediction but also other alternative hypotheses, as mentioned in [6, 60]. This approach refers to hypothesis-driven decision-making, which can help counter automation bias and fixation by encouraging users to consider evidence of multiple hypotheses. Furthermore, Rastogi et al. [70] proposed a time-based strategy to address anchoring bias by allocating time according to AI confidence.
2.3 Decision Support Paradigms
@@ 2.3.1 AI-first Workflow
2.3.2 Human-first Workflow
Human-first decision workflow has been shown to help reduce the over-reliance on erroneous AI recommendations [9, 21]. However, experts may interact with decision-making systems differently from laypeople (crowdworkers). For example, Fogliato et al. [21], Gaube et al. [24] ran studies with radiologists who were the experts. Their task is to review patients’ X-ray images. The studies conclude that in human-first workflow with expert participants, they are less likely to leverage AI advice even though the AI is more accurate. In fact, this is referred to algorithm aversion [20] or under-reliance.
A human-first approach called cognitive forcing, based on earlier ideas in psychology for interventions that elicit human thinking at decision-making time [48], has been proposed as a way to improve users’ engagement and also increase their learning when interacting with the AI [9, 23]. Four cognitive forcing designs have been introduced: (1) On demand: Participants can only see the AI recommendation when they request it; (2) Update: Participants first made a decision without seeing the AI recommendation. Then, they were shown the AI prediction and could update their decision later; (3) Wait: Participants had to wait for 30 seconds before the AI decision was shown; and (4) Only AI explanation: Providing just the AI explanation and no AI recommendation, on the basis that this may help people process the AI explanation more carefully and therefore, improve their knowledge and make better decisions [23]. Importantly, cognitive forcing has been shown to reduce over-reliance compared to the standard AI suggestion approach, although that study had a limitation in that the AI prediction was always correct. There are also some recognised trade-offs of cognitive forcing designs: more time-consuming [23], and less trust [9].
2.3.3 Evaluative AI (Hypothesis-driven) Paradigm
Miller [60] argues that AI-assisted decision support is on the cusp of a paradigm shift. This shift is away from the idea of human-first or AI-first, and into a framework he calls evaluative AI, which is built around the Data/Frame model [41]. The key insight of evaluative AI is to not necessarily provide a recommendation, and instead to support the human cognitive decision-making process by providing evidence for or against the particular hypothesis that a human decision-maker is considering. This would help to prevent over- and under-reliance, and would help the decision maker to retain their internal locus of control [74].
@@ 2.4 Explainable AI (XAI) in Decision Support
In this section, we will review some AI-first explainable AI approaches that have been used to provide explanations for the AI recommendation based on feature analysis. We will also discuss some popular evidence-based explanations that have been used in supporting decision-making.
2.4.1 Human Decision-Making Reliance on AI Support
@@
There is no straightforward position regarding when humans are well-calibrated to accept AI-generated advice [82]. Overall, study participants appear more likely to accept an AI’s recommendation when provided with explanations, regardless of the model’s correctness [5, 30, 83, 10], whereas less detailed explanations can lead to self-reliance [10]. Explanations can increase the accuracy of the human-AI team when the AI is correct, but decrease it when it is wrong, resulting in over-reliance on the AI’s recommendations. Arguably, this is because the current explanation forms do not provide details of the underlying rationale of the AI model behaviour [83]. We therefore should be careful when selecting the explanation type as it can have a significant effect on whether users decide to rely on them  [13, 55]. Moreover, Vasconcelos et al. [81] found that explanations can reduce over-reliance by increasing the task difficulty, easing the explanation difficulty or increasing the benefit of task completion via monetary rewards.
2.4.2 Explanations and Cognitive Engagement
Besides providing AI recommendations, we now review alternative approaches that can help improve human decision-making. Gajos and Mamykina [23] suggested that providing just the AI explanation and no AI recommendation can help people process the AI explanation more carefully and therefore, improve their knowledge and make better decisions. Cheng et al. [16] showed that interactive interfaces can improve users’ understanding of how the AI algorithm works and reduce over-reliance on the AI with cognitive forcing functions [9]. However, cognitive forcing functions can only improve the performance of human+AI teams compared to only explanations given in situations when the AI prediction is incorrect. Naiseh et al. [63] investigated how friction – small design elements that slow users down – can improve trust calibration in XAI systems by nudging users to engage more thoughtfully with AI-generated explanations. A study with medical practitioners showed increased interaction with explanations under friction conditions, though it did not significantly deepen cognitive engagement.
Along the same lines as our work, Cabitza et al. [11] introduced a reflective XAI support that makes users reflect on the reliability of the AI advice by providing evidence both in favour of and against the recommendation. The evidence was presented by showing similar cases, but one where the AI was correct and one where it failed. While the results showed that this approach did not improve diagnostic accuracy, it was appreciated by experienced users for promoting thoughtful evaluation. Cabitza et al. [12] then explored a form of explainable AI (XAI) called pro-hoc explanations, which present users with similar example cases for each possible outcome instead of offering a single AI prediction. While it led to only modest accuracy improvements, it significantly increased user confidence, especially among less experienced clinicians. Ma et al. [57] proposed Deliberative AI, built on the Weight of Evidence (WoE) framework [59], to provide evidence for conflicting opinions. The results found that Deliberative AI can foster appropriate human reliance on AI and improve decision-making performance.
Unlike the above approaches, we not only use existing methods but also introduce a novel method. To our knowledge, this is the first work that proposes a general tool for evaluative AI across both tabular and image datasets. We also conducted two user studies – one with large-scale crowdsourcing and another with a small group of experienced participants – to empirically evaluate the effectiveness of our approach in improving human decision-making.
2.4.3 Evidence-based Explanations
We describe key differences between the Weight of Evidence approach (WoE) and other feature attribution approaches such as LIME [72] and SHAP [56].
@@ 2.4.3 Evidence-based Explanations
An approach to generate evidence is the Weight of Evidence (WoE) framework [59], which can be adapted to meet human-centred design principles. Formally, given a predicted output $y$ for an input $x$, WoE seeks how much evidence the input feature $x_{i}$ gives in favour of (or against) $y$. WoE is similar to feature importance explanations. However, the main difference is that Weight of Evidence uses log likelihoods and log odds ratios to generate explanations, whereas LIME [72] and SHAP [56] find feature importance by modifying the predictive posterior probability in various ways.
We choose the Weight of Evidence approach (WoE) because WoE follows human-centred design principles [59], in which explanations should be contrastive (i.e., why the model predicted $y$ instead of alternative $y^{\prime}$), exhaustive (i.e., justify on why every alternative $y^{\prime}$), compositional (i.e., be able to break down into simple components in the prediction), easily-understandable (i.e., understandable components) and parsimonious (i.e., only provide most relevant facts). Kumar et al. [46] argue that SHAP [56] has several human-centred issues, including non-contrastive and non-actionable explanations, and that most people do not have a correct mental model of Shapley values. LIME [72] does not follow human-centred design principles either. Moreover, LIME [72] uses a surrogate model to approximate the original one, leading to lower performance and explanations that do not reflect the original model. SHAP [56] can have a high computational cost due to the need to compute Shapley values for each instance. In contrast, WoE [59] is a probabilistic approach that does not require a surrogate model and can be computed more efficiently.
@@
A closely related work to Melis et al. [59] is from Poulin et al. [69]. Poulin et al. [69] propose a framework called ExplainD that uses additive evidence. The framework also measures the weight of evidence using a Naive Bayes classifier along with highlighting the negative and positive evidence for a decision. However, the problem being considered is a binary classification. Furthermore, there is still room for improvement by conducting experiments to evaluate the framework. Kulesza et al. [45, 44] introduce EluciDebug in email classification using Multinomial Naive Bayes classifier (MNB). The EluciDebug prototype provides an interface that includes important words and the folder size that both contribute to the email classification. In this example, important words can be referred to as strength of evidence. The folder size which describes the number of instances in each class can be referred to weight of evidence. Yet the prototype did not specifically give positive and negative evidence in decision-making situations.
Having more evidence does not always have positive effects on decision-making. Ratcliff and Smith [71] propose stopping rules that can be applied when additional evidence does not change the final decision. More specifically, the relative stopping rule refers to when the balance between hypotheses reaches a threshold, meaning evidence for one alternative inherently counts against the other. On the other hand, the absolute stopping rule leads to a decision when evidence for a single hypothesis reaches a threshold, with alternatives being considered independently.
3 Methodology
In this section, we present the methodology for our proposed Evaluative AI model, We first introduce the Weight of Evidence (WoE) framework by defining the concept of evidence for both independent and dependent variables. Since tabular data provides pre-defined features, we can directly apply the WoE framework. Then, we introduce the Visual Evaluative AI model, which combines concept-based explanations with the WoE framework, for generating evidence from image data.
3.1 Weight of Evidence